How do I find the value of y so that the line goes the through the points (2,-2) and (x,3) and has a slope of 5. I already did the y=(y2-y1) over (x2-x1)I don't know what to do now. Please help!!!
Answer by TimothyLamb(4379)

---
given:
(2,-2)
slope = 5
---
to solve the linear equation, copy and paste this (the point of intersection and the parallel slope):
2,-2,5
into the "Point-Slope form: x1 y1 m" input box here: https://sooeet.com/math/linear-equation-solver.php
---
x-intercept = (2.4, 0)
y-intercept = (0, -12)
slope = 5
---
answer:
the slope-intercept form of our new line:
y = 5x - 12
---
solving manually:
---
the general point-slope form:
y - y1 = m(x - x1)
---
given:
(2,-2)
slope = 5
y + 2 = 5(x - 2)
y = 5x - 10 - 2
---
y = 5x - 12
---
NOTE:
the linear equation solver and the manual solution agree.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php